OpenStudy (anonymous):

Verify using sum &difference identities: sin(x+y)-sin(x-y)=2cosxsiny

OpenStudy (alexwee123):

sin(x+y)-sin(x-y) sinxcosy+cosxsiny-sinxcosy+cosxsiny sinxcosy -sinxcos+cosxsiny+cosxsiny =2cosxsiny

OpenStudy (callisto):

Just to provide another method: sinA-sinB = 2cos[(A+B)/2]sin[(A-B)/2] Put A=x+y B=x-y into the above formula sin(x+y)-sin(x-y) = 2cos{(x+y)+(x-y)]/2}sin{(x+y)-(x-y)]/2} = 2 cos (2x/2) sin (2y/2) = 2cosxsiny

OpenStudy (anonymous):

@alexwee123 what did u do on the third step?

OpenStudy (alexwee123):

i arranged the values so it can be clear that sinxcosy -sinxcosy would be cancelled out

OpenStudy (anonymous):

ok thx ^^